OpenStudy (anonymous):

670mL of 2M soduim chloride (NaCl)

OpenStudy (anonymous):

Yes..?

OpenStudy (anonymous):

im tryna learn how to work it out

OpenStudy (anonymous):

C1*V1=C2*V2 C= Concentration V= Volume. You're lacking one variable to solve this equation, you need to know either what concentration C2 or V2 has, If these are your values for C1 & V1. If its your C2 or V2 (desired concetration & volume) You need to figure out either C1 or V1 inorder to solve this. You should visist the Chemistry group for this type of questions in the future.

OpenStudy (anonymous):

now thst i have my fw how can i break down 670ml

OpenStudy (anonymous):

Decreasing V therefore increasing C. Another hint is "which has lowest boiling point Water or NaCl?"

OpenStudy (blues):

The chemistry group is found here: http://openstudy.com/study#/groups/Chemistry.